# Unleashing the Potential of Large Language Models as Prompt Optimizers: An Analogical Analysis with Gradient-based Model Optimizers

### v1

@@ 1 Introduction
Nowadays, prompting has become the pivotal approach to unleashing the power of large language models (LLMs) Zhao et al. (2023); Touvron et al. (2023); OpenAI (2023). However, prompt engineering is not easy and requires extensive trial-and-error efforts since LLMs are sensitive to prompts Zhao et al. (2021); Lu et al. (2022); Wei et al. (2023). Although general guidelines for high-quality prompts exist Kojima et al. (2022); Amatriain (2024), they cannot always lead to optimal task performance.
To improve the task performance of LLMs, automatic prompt optimization has been proposed Zhou et al. (2023). Early work either directly performs discrete optimization through methods like reinforcement learning Deng et al. (2022); Zhang et al. (2023) or performs continuous optimization in the embedding space of LLMs Shin et al. (2020); Lin et al. (2023); Chen et al. (2023). However, these methods often require access to the logits or internal states of LLMs, which is infeasible for those only accessible through APIs. In addition, they need to be specially trained for each task. Considering these issues, recent work proposes to model the optimization problem in natural language and using LLMs as prompt optimizers Zhou et al. (2023); Yang et al. (2023); Ma et al. (2024). In this approach, LLMs can perform optimization with only outputs from the task model and quickly adapt to various tasks without training. However, such a method raises a new challenge for the design of meta-prompt, which is the prompt for LLMs to perform prompt optimization.
@@
Specifically, in this paper, we take the gradient-based optimization approach from machine learning as reference, and identify two key factors for model parameter learning, namely update direction (i.e., gradient calculation) and update method (i.e., gradient descent). By analogy, in LLM-based prompt optimizers, the update direction can refer to suitable prompt information informing the LLM about how to produce more effective task prompts, while the update method can refer to how the LLM utilizes such information to improve prompts. By drawing such an analogy, we can develop a more formal framework for LLM-based prompt optimizers and provide guidelines to design more principled meta-prompts.
@@
2 Related Work
Our work is related to the following two directions.
@@ 3.2.1 Revisiting Gradient-Based Optimizers
Factor
Gradient-based
model optimizer
LLM-based
prompt optimizer
Update direction Model value Prompt
Gradient Reflection
Momentum Trajectory
Update method Learning rate Edit distance
Descent Editing / Generation
@@ 3.2.3 Analogical Prompt Optimization Strategies for “Update Method
Prompt Variation Control. After setting the optimization direction, it is important to adjust the parameter optimization process with a suitable learning rate. This is because using a either too large or too small rate can result in an oscillating or slowly converging optimization process. Similarly, without a suitable control mechanism in prompt optimization, the LLM-based optimizer might overshoot the optimal prompt or oscillate in the optimization process. To mimic the similar effects of learning rate, our idea is to control the variation degree of prompt optimization, which is measured by the edit distance between two task prompts at consecutive iterations. Specifically, following Ye et al. (2023), we add an instruction (i.e., “You are allowed to change up to X words in the current prompt.”) into the meta-prompt to limit the number of words that can be modified. Accordingly, we propose two control methods as follows:
@@
$\bullet$ Editing-based refinement. The first method directly edits the last-round task prompt to improve performance. Specifically, we add an instruction (i.e., “Modify the current prompt and get a new improved promp”) into the meta-prompt, which requires the LLM to edit the task prompt from the last step according to the update direction. This method allows for effectively exploiting the existing prompt, leading to a gradual performance improvement through a relatively stable optimization process.
[…]
@@ 3.3.1 Experiment Setup
We select a dataset from each type of task in BBH Suzgun et al. (2023) to create a lite BBH benchmark for our analysis: i) Navigate (binary choice); ii) Movie Recommendation (multiple choice); iii) Object Counting (numeric response); iv) Word Sorting (free response). We adopt exact match as the metric for performance evaluation. We employ Llama-2-7b-chat Touvron et al. (2023) as the task model and gpt-3.5-turbo as the prompt optimizer. The initial task prompt is “Let’s think step by step.” Other details are presented in Appendix B.
@@ 3.3.2 Empirical Findings
$\bullet$ In general, generation-based refinement performs better than editing-based refinement, which brings an improvement of up to 36% compared with the initial task prompt. This may be because LLMs are not specially trained for prompt optimization, and the demonstrations in the generation-based strategy can provide guidance about this task, thus helping LLMs learn to perform better.
@@ 4 Our Prompt Optimizer: GPO
$\bullet$ Update direction. To determine the update direction, we leverage the retrieval-based optimization trajectory in Section 3.2.2. This trajectory comprises a collection of past task prompts, along with their model performance. They are selected using a relevance-based strategy and are sorted in ascending order based on their performance.
@@
Furthermore, we enhance the meta-prompt by incorporating a few task examples. These examples provide additional context to aid the LLM in understanding the task more effectively. The complete meta-prompt is presented in Appendix C.2.
@@ 5.1 Experimental Setup
Tasks and Datasets. We select datasets from three groups of tasks for the experiment: Big-Bench Hard (BBH) Suzgun et al. (2023) and GSM8K Cobbe et al. (2021) for complex reasoning tasks, MMLU Hendrycks et al. (2021) for knowledge-intensive tasks, and WSC Levesque et al. (2012) and WebNLG Gardent et al. (2017) for common NLP tasks. Due to computational limitations, we sample a subset from each dataset for the main experiment. In addition, we use the lite BBH benchmark in Section 3.3.1 for detailed analysis. Other details are presented in Appendix C.1.
@@
Implementation Details. For the task model, we use both the base model (i.e., Llama-2-7b) and the instruction-tuned models (i.e., Llama-2-7b-chat, GPT-3.5-turbo, and GPT-4). For the prompt optimizer, we use GPT-3.5-turbo and GPT-4. Unless otherwise specified, we use Llama-2-7b-chat as the task model and GPT-3.5-turbo as the prompt optimizer throughout experiments. For the initial task prompt, we use the original ones from Hendrycks et al. (2021) for MMLU, “Let’s think step by step.” from Kojima et al. (2022) for GSM8K and BBH, and “Let’s solve the problem.” from Yang et al. (2023) for WSC and WebNLG. We repeat all the experiments three times and report the average of the results. Other details are presented in Appendix C.2.
@@ 5.3 Detailed Analysis
The Impact of Initial Prompts.
@@
In our main experiment, we take “Let’s think step by step.” as the initial prompt. In this part, we aim to explore the impact of initial task prompts. Specifically, we consider prompts from three categories (instructive, misleading, and irrelevant) and select three prompts from each category following Kojima et al. (2022). Table 8 presents the results on the lite BBH benchmark. In general, GPO can consistently boost performance across various initial prompts. This observation underscores the versatility of GPO as a prompt optimizer. Furthermore, the efficacy of optimization is more pronounced with relevant initial prompts (i.e., instructive and misleading). It means that prompt initialization is still important, especially in conveying task-specific information.
The Efficiency of Optimization.
(a) Optimization curve
(b) Token consumption
LLM-based prompt optimization requires multiple interactions with the LLM. In this part, we investigate the efficiency of GPO from the optimization curve and token consumption. First, Figure 0(a) shows that on the movie recommendation dataset, GPO exhibits rapid enhancement of the task prompt in the early stage, followed by consistent performance improvement in the later stage of optimization. Second, as depicted in Figure 0(b), the average token consumption of GPO on the lite BBH benchmark is much lower than SGDM, APO, and PE2, and comparable to APR and OPRO. Since GPO only utilizes task prompts to derive the update direction and performs fine-grained control over the variation, it can achieve better performance with high efficiency.
6 Conclusion
We present GPO, a novel gradient-inspired LLM-based prompt optimizer. It utilizes LLMs to automatically optimize prompts, drawing inspiration from gradient-based model optimization techniques. By identifying the two crucial aspects of update direction and update method in model optimization, we enhance prompt optimizers by incorporating concepts from gradient-based model optimizers. Through extensive experiments, GPO demonstrates remarkable capabilities for prompt optimization across diverse tasks, models, and initial prompts. Moreover, it surpasses competitive baselines while consuming fewer tokens.
7 Limitations
Appendix A Hyper-Parameter Analysis
@@
Appendix B Additional Details for the Setup of Analogical Analysis
@@
Implementation Details. Following Crispino et al. (2023), we select Llama-2-7b-chat Touvron et al. (2023) as the task model and set its temperature to 0 to make the output as deterministic as possible. For the prompt optimizer, we employ gpt-3.5-turbo, which is the underlying model of ChatGPT. We set its temperature to 1.0 to encourage the generation to be more diverse. To help the prompt optimizer understand the downstream task, following Yang et al. (2023), we randomly sample 3 examples from the dataset and fill them into the meta-prompt of the prompt optimizer. In the process of optimization, we take the instruction “Let’s think step by step.” as the initial prompt and insert it at the end of the question to obtain better performance following Suzgun et al. (2023). At each step, we first prompt the optimizer to generate 8 candidate task prompts, and then select the best-performing one as the task prompt for the next iteration following Yang et al. (2023); Ye et al. (2023). The optimization process lasts for at most 3 epochs with a batch size of 8. We repeat all the experiments three times and report the average of the results. The meta-prompts we used are detailed in Appendix D.1.
Appendix C Additional Details for the Setup of Experiment
C.1 The Statistics of Datasets
@@
C.2 Implementation Details
For the task model, we use both the base model (i.e., Llama-2-7b) and the instruction-tuned models (i.e., Llama-2-7b-chat, GPT-3.5-turbo, and GPT-4). The temperature is 0. For the prompt optimizer, we use GPT-3.5-turbo and GPT-4. The temperature is 1.0. Unless otherwise specified, we use Llama-2-7b-chat as the task model and GPT-3.5-turbo as the prompt optimizer throughout experiments. In the meta-prompt, we include 3 task examples and 7 historical task prompts. For the initial task prompt, we use the original ones from Hendrycks et al. (2021) for MMLU, “Let’s think step by step.” from Kojima et al. (2022) for GSM8K and BBH, and “Let’s solve the problem.” from Yang et al. (2023) for WSC and WebNLG. At each step, the optimizer generates 8 candidates, and the best-performing one is selected. The optimization lasts for at most 3 epochs with a batch size of 8. We repeat all the experiments three times and report the average of the results. The meta-prompts are detailed in Appendix D.2.
Appendix D Meta-Prompt
D.1 Analogical Analysis

### v2

@@ 1 Introduction
Factor
Gradient-based
model optimizer
LLM-based
prompt optimizer
Update direction Model value Task prompt
Gradient Reflection
Momentum Trajectory
Update method Learning rate Edit distance
Descent Refinement strategy
Figure 1: The overview of LLM-based prompt optimization framework.
Nowadays, prompting has become the pivotal approach to unleashing the power of large language models (LLMs) Touvron et al. (2023); OpenAI (2023). However, prompt engineering is not easy and requires extensive trial-and-error efforts since LLMs are sensitive to prompts Zhao et al. (2021); Lu et al. (2022); Wei et al. (2023). Although general guidelines for high-quality prompts exist Kojima et al. (2022); Amatriain (2024), they cannot always lead to optimal task performance.
To improve the task performance of LLMs, automatic prompt optimization has been proposed Zhou et al. (2023). Early work either directly performs discrete optimization through methods like reinforcement learning Deng et al. (2022); Zhang et al. (2023) or performs continuous optimization in the embedding space of LLMs Shin et al. (2020); Lin et al. (2023); Chen et al. (2023). However, these methods often require access to the logits or internal states of LLMs, which is infeasible for those only accessible through APIs. In addition, they need to be specially trained for each task. Considering these issues, recent work proposes to model the optimization problem in natural language and using LLMs as prompt optimizers Zhou et al. (2023); Ma et al. (2024); Schnabel and Neville (2024). In this approach, LLMs can perform optimization with only outputs from the task model and quickly adapt to various tasks without training. Figure 1 illustrates the overall framework of LLM-based prompt optimization. However, such a method raises a new challenge for the design of meta-prompt, which is the prompt for LLMs to perform prompt optimization.
@@
Specifically, in this paper, we take the gradient-based optimization approach from machine learning as reference, and identify two key factors for model parameter learning, namely update direction (i.e., gradient calculation) and update method (i.e., gradient descent), as presented in Table 1. By analogy, in LLM-based prompt optimizers, the update direction can refer to suitable prompt information informing the LLM about how to produce more effective task prompts, while the update method can refer to how the LLM utilizes such information to improve prompts. By drawing such an analogy, we can develop a more formal framework for LLM-based prompt optimizers and provide guidelines to design more principled meta-prompts.
@@ 2.2.3 Analogical Prompt Optimization Strategies for “Update Method
Prompt Variation Control. After setting the optimization direction, it is important to adjust the parameter optimization process with a suitable learning rate. This is because using a either too large or too small rate can result in an oscillating or slowly converging optimization process. Similarly, without a suitable control mechanism in prompt optimization, the LLM-based optimizer might overshoot the optimal prompt or oscillate in the optimization process. To mimic the similar effects of learning rate, our idea is to control the variation degree of prompt optimization, which is measured by the edit distance between two task prompts at consecutive iterations. Specifically, following Ye et al. (2023), we add an instruction (i.e., “You are allowed to change up to {number} words in the current prompt.”) into the meta-prompt to limit the number of words that can be modified. Accordingly, we propose two control methods as follows:
@@
$\bullet$ Editing-based refinement. The first method directly edits the last-round task prompt to improve performance Pryzant et al. (2023). Specifically, we add an instruction (i.e., “Modify the current prompt and get a new improved promp”) into the meta-prompt, which requires the LLM to edit the task prompt from the last step according to the update direction. This method allows for effectively exploiting the existing prompt, leading to a gradual performance improvement through a relatively stable optimization process.
$\bullet$ Generation-based refinement. In addition to direct edits, we can also leverage the in-context learning capability of LLMs to generate refined task prompts Zhou et al. (2023). Specifically, we present the information regarding the update direction in the format of demonstration. Then, the LLM can follow the demonstration to generate a new task prompt. Compared with the editing-based updating method, the generation-based approach explores a wider range of prompt variations, which has the potential to yield improved performance but also makes the optimization process unstable.
@@ 2.3.1 Experiment Setup
We select a dataset from each type of task in BBH Suzgun et al. (2023) to create a lite BBH benchmark for our analysis: i) Navigate (binary choice); ii) Movie Recommendation (multiple choice); iii) Object Counting (numeric response); iv) Word Sorting (free response). We adopt exact match as the metric for performance evaluation. We employ Llama-2-7b-chat Touvron et al. (2023) as the task model and gpt-3.5-turbo as the prompt optimizer. The initial task prompt is “Let’s think step by step.” Other details are presented in Appendix C.
@@ 2.3.2 Empirical Findings
$\bullet$ In general, generation-based refinement performs better than editing-based refinement, which brings an improvement of up to 36% compared with the initial task prompt. This can be partly attributed to that exploration is more important in prompt optimization. The generation-based strategy is not confined to the current prompt, allowing LLMs to explore a boarder search space, thereby demonstrating greater flexibility and creativity.
@@ 3 Our Prompt Optimizer: GPO
$\bullet$ Update direction. To determine the update direction, we leverage the retrieval-based optimization trajectory in Section 2.2.2. This trajectory comprises a collection of past task prompts, along with their model performance. They are selected using a relevance-based strategy and are sorted in ascending order based on their similarity to the current prompt.
@@
Furthermore, we enhance the meta-prompt by incorporating a few task examples. These examples provide additional context to aid the LLM in understanding the task more effectively. The complete meta-prompt is presented in Appendix D.2.
@@ 4.1 Experimental Setup
Tasks and Datasets. We select datasets from three groups of tasks for the experiment: Big-Bench Hard (BBH) Suzgun et al. (2023) and GSM8K Cobbe et al. (2021) for complex reasoning tasks, MMLU Hendrycks et al. (2021) for knowledge-intensive tasks, and WSC Levesque et al. (2012) and WebNLG Gardent et al. (2017) for common NLP tasks. Due to computational limitations, we sample a subset from each dataset for the main experiment. In addition, we use the lite BBH benchmark in Section 2.3.1 for detailed analysis. Other details are presented in Appendix D.1.
@@
Implementation Details. For the task model, we use both the base model (i.e., Llama-2-7b) and the instruction-tuned models (i.e., Llama-2-7b-chat, GPT-3.5-turbo, and GPT-4). For the prompt optimizer, we use GPT-3.5-turbo and GPT-4. Unless otherwise specified, we use Llama-2-7b-chat as the task model and GPT-3.5-turbo as the prompt optimizer throughout experiments. We repeat all the experiments three times and report the average of the results. Other details are presented in Appendix D.2.
@@ 4.3 Detailed Analysis
The Efficiency of Optimization.
(a) Optimization curve
(b) Token consumption
LLM-based prompt optimization requires multiple interactions with the LLM. In this part, we investigate the efficiency of GPO from the optimization curve of the first 12 steps and token consumption. First, Figure 2(a) shows that on the movie recommendation dataset, GPO exhibits rapid enhancement of the task prompt in the early stage, followed by consistent performance improvement in the later stage of optimization. Second, as depicted in Figure 2(b), the average token consumption of GPO on the lite BBH benchmark is much lower than SGDM, APO, and PE2, and comparable to APE and OPRO. Since GPO only utilizes task prompts to derive the update direction and performs fine-grained control over the variation, it can achieve better performance with high efficiency.
5 Related Work
Our work is related to the following two directions.
6 Conclusion
We present GPO, a novel gradient-inspired LLM-based prompt optimizer. It utilizes LLMs to automatically optimize prompts, drawing inspiration from gradient-based model optimization techniques. By identifying the two crucial aspects of update direction and update method in model optimization, we enhance prompt optimizers by incorporating concepts from gradient-based model optimizers. Through extensive experiments, GPO demonstrates remarkable capabilities for prompt optimization across diverse tasks, models, and evaluation settings. Moreover, it surpasses competitive baselines while consuming fewer tokens.
7 Limitations
Appendix A Extended Analysis
A.1 The Convergence of LLM-based Prompt Optimization
Figure 3: Optimization curve on Navigate dataset.
Figure 4: Optimization curve on Movie Recommendation dataset.
Figure 5: Optimization curve on Object Counting dataset.
Figure 6: Optimization curve on Word Sorting dataset.
The convergence of LLM-based prompt optimization occurs when performance stabilizes at a plateau during the iterative process. Here, we illustrate the entire optimization curve on the lite BBH benchmark in Figure 3, Figure 4, Figure 5, and Figure 6. It can be observed that all the prompt optimization methods tend to reach a stable state within 12 steps. Therefore, our experimental settings (i.e., 3 epochs) are sufficient to demonstrate the effectiveness of GPO.
A.2 The Impact of Initial Prompts
@@
In our main experiment, we take “Let’s think step by step.” as the initial prompt. In this part, we aim to explore the impact of initial task prompts. Specifically, we consider prompts from four categories (instructive, misleading, irrelevant, and empty) following Kojima et al. (2022). Table 8 presents the results on the lite BBH benchmark. In general, GPO can consistently boost performance across various initial prompts. This observation underscores the versatility of GPO as a prompt optimizer. Furthermore, the efficacy of optimization is more pronounced with relevant initial prompts (i.e., instructive and misleading). It means that prompt initialization is still important, especially in conveying task-specific information.
A.3 The Generalizability of Optimized Prompts
Method Prompt Accuracy
MultiArith AQUA
Zero-shot CoT Let’s think step by step. 58.00 19.69
OPRO
Take a deep breath and work on this problem step-by-step.
59.33 20.47
GPO
Thoroughly strategize and meticulously plan every single
aspect of your approach, leaving no room for oversight,
and give utmost attention to even the smallest details,
with the goal of achieving exceptional results and excelling.
62.33 24.80
Table 9: Generalizability across datasets: Accuracies of optimal prompts for GSM8k on MultiArith and AQUA.
To assess the generalizability of optimized prompts, following Zhou et al. (2023), we evaluate the prompts optimized on GSM8k Cobbe et al. (2021) by OPRO and GPO on two additional math reasoning benchmarks MultiArith Roy and Roth (2016) and AQUA Ling et al. (2017). Table 9 demonstrates that the optimized prompts by GPO can be generalized well on these two benchmarks, which perform better than those by OPRO.
A.4 Hyper-Parameter Analysis
@@
A.5 The Execution Time of LLM-based Prompt Optimizers
Figure 8: Execution time of LLM-based prompt optimizers
To further assess the computation costs of LLM-based prompt optimizers, we conduct five repetition experiments and record their average execution time. Figure 8 shows the results on the lite BBH benchmark. We can observe that the execution time is almost consistent with the trend of token consumption (in Section 7). The average execution time of GPO is much lower than SGDM, APO, and PE2, and comparable to OPRO.
Appendix B The Interpretability of Optimized Prompts
Step Prompt
0
Let’s think step by step.
1
Plan your steps carefully.
2
Strategize your approach meticulously.
3
Strategically plan your approach down to the smallest details.
4
Carefully strategize and meticulously plan your approach down to the tiniest details.
5
Thoroughly strategize and meticulously plan every aspect of your approach,
ensuring attention to even the most minuscule details.
6
Strive to meticulously strategize and thoroughly plan each and every aspect of your approach,
leaving no room for oversight, and paying utmost attention to even the most infinitesimal details.
9
Thoroughly strategize and meticulously plan every single aspect of your approach, ensuring there is no room
for oversight, and paying utmost attention to even the minutest details, in order to achieve perfect results.
12
Thoroughly strategize and meticulously plan every single aspect of your approach, leaving no room for oversight,
and give utmost attention to even the smallest details, with the goal of achieving exceptional results and excelling.
Table 10: Intermediate prompts optimized by GPO on GSM8k dataset.
Step Prompt
0
Let’s think step by step.
1
Thinking step by step will help us find the solution.
2
Sort the given list of words in alphabetical order.
3
Arrange the given list of words in ascending order based on their alphabetical order.
4
Arrange the given list of words in alphabetical order, from A to Z.
5
Sort the given words in alphabetical order, from A to Z.
6
Arrange the given words in alphabetical order, from A to Z.
9
Arrange the given list of words in ascending alphabetical order, starting from the word with the
lowest letter and ending with the word with the highest letter, disregarding case sensitivity.
12
Sort the given list of words in ascending alphabetical order, starting from the word with the lowest letter
and ending with the word with the highest letter, while considering both uppercase and lowercase letters.
Table 11: Intermediate prompts optimized by GPO on Word Sorting dataset.
To further enhance the interpretability of optimized prompts, we present the intermediate prompts optimized by GPO on GSM8k and Word Sorting dataset in Table 10 and Table 11. It can be observed that as the number of steps increases, the optimized prompts contain more task-related information (e.g., “Sort the given list of words in ascending alphabetical order, starting from the word with the lowest letter and ending with the word with the highest letter”) and fine-grained hints (e.g., “Thoroughly strategize and meticulously plan every single aspect of your approach, leaving no room for oversight, and give utmost attention to even the smallest detail”) to better solve the task. In addition, we find that the optimized prompts do not contain information on specific task samples, leading to better generalizability and robustness.
Appendix C Additional Details for the Setup of Analogical Analysis
@@
Implementation Details. Following Crispino et al. (2023), we select Llama-2-7b-chat Touvron et al. (2023) as the task model and set its temperature to 0 to make the output as deterministic as possible. For the prompt optimizer, we employ gpt-3.5-turbo, which is the underlying model of ChatGPT. We set its temperature to 1.0 to encourage the generation to be more diverse. To help the prompt optimizer understand the downstream task, following Yang et al. (2023), we randomly sample 3 examples from the dataset and fill them into the meta-prompt of the prompt optimizer. In the process of optimization, we take the instruction “Let’s think step by step.” as the initial prompt and insert it at the end of the question to obtain better performance following Suzgun et al. (2023). At each step, we first prompt the optimizer to generate 8 candidate task prompts, and then select the best-performing one as the task prompt for the next iteration following Yang et al. (2023); Ye et al. (2023). The optimization process lasts for at most 3 epochs with a batch size of 8. We repeat all the experiments three times and report the average of the results. The meta-prompts we used are detailed in Appendix E.1.
Appendix D Additional Details for the Setup of Experiment
D.1 The Statistics of Datasets
@@
D.2 Implementation Details
For the task model, we use both the base model (i.e., Llama-2-7b) and the instruction-tuned models (i.e., Llama-2-7b-chat, GPT-3.5-turbo, and GPT-4). The temperature is 0. For the prompt optimizer, we use GPT-3.5-turbo and GPT-4. The temperature is 1.0. Unless otherwise specified, we use Llama-2-7b-chat as the task model and GPT-3.5-turbo as the prompt optimizer throughout experiments. In the meta-prompt, we include 3 task examples and 7 historical task prompts. For the initial task prompt, we use the original ones from Hendrycks et al. (2021) for MMLU, “Let’s think step by step.” from Kojima et al. (2022) for GSM8K and BBH, and “Let’s solve the problem.” from Yang et al. (2023) for WSC and WebNLG. At each step, the optimizer generates 8 candidates, and the best-performing one is selected. The optimization lasts for at most 3 epochs with a batch size of 8. We repeat all the experiments three times and report the average of the results. The meta-prompts are detailed in Appendix E.2.
Appendix E Meta-Prompt
E.1 Analogical Analysis